The highest Score in an innings was 3/11 of the total and the next highest was 3/11 of the reminder. If the scores differed by 9, find the total score.

Grade:

1 Answers

divanshu anand
12 years ago

Dear Hemant

Let the total score be X then first highest = 3/11 x

then next highest is 3/11th of the remainder ie 3/11 * (1-3/11)

subtracting the two we get 3/11 - 3/11(1-3/11)

3/11(1-8/11)x = 9

3/11(3/11)x = 9

x = 9 * 11 * 11 / 3 *3 = 121

so x= 121

ie total runs =121 ans...